OpenStudy (anonymous):

Simplify the expression radical -4 over (5+2i)-(3-4i) :))

OpenStudy (dumbcow):

what is inside the radical?

OpenStudy (anonymous):

-4 :)

OpenStudy (dumbcow):

sqrt(-4) = 2i multiply by the conjugate on top and bottom to get rid of i in the denominator

OpenStudy (anonymous):

wait would i be able to FOIL for this equation?

OpenStudy (dumbcow):

yes you will have to

OpenStudy (dumbcow):

\[\frac{2i}{2+6i} = \frac{i}{1+3i}\] \[\frac{i}{1+3i}*\frac{1-3i}{1-3i}\]

OpenStudy (anonymous):

okay thank you:)